# Dropping Just a Handful of Preferences Can Change Top Large Language Model Rankings

### v2

@@ 5.2 Data-dropping Robustness
A growing body of works in statistics and theoretical computer science develops algorithms for assessing whether data analyses are robust to dropping a small, worst-case fraction of the data (Broderick et al., 2020; Kuschnig et al., 2021; Moitra and Rohatgi, 2023; Freund and Hopkins, 2023; Nguyen et al., 2024; Huang et al., 2025a; Rubinstein and Hopkins, 2025). To our knowledge, only one prior work has investigated this question in the context of ranking systems: Shiffman et al. (2023) study the robustness of rankings in gene set enrichment analysis, showing that dropping just a few cells can alter the ranking of p-values derived from the hypergeometric test. In contrast, our work examines ranking robustness in a BT-based ranking system. While Shiffman et al. (2023) analyze p-value rankings, we analyze preference-based rankings of LLMs, extending approximation methods such as AMIP (Broderick et al., 2020) and Additive One-step Newton (Huang et al., 2025a) to study the robustness of BT-based ranking systems.
@@ 6 Discussion
Sensitivity to worst-case data-dropping is often indicative of low signal-to-noise in the underlying data (Broderick et al., 2020); to help increase signal-to-noise, we recommend three different design-related improvements that AI arenas could take. (1) Collect richer forms of feedback beyond binary preferences (e.g., asking for evaluators’ confidence levels (Méndez et al., 2022). [^6] The weighted logistic regression model used by Chatbot Arena can easily be extended to take in confidence ratings on top of binary preferences.One could imagine implementing this through encoding the confidence rating as a weight in the Win-Counts matrix described in the “Chatbot Arena Leaderboard Calculation (Bradley–Terry model)” Colab notebook. (2) Design more discriminative prompts. Arenas could incorporate a prompt-filtering system to identify and remove uninformative prompts, or create tools to identify prompts requiring specialized knowledge in order to route them to appropriate evaluators (Don-Yehiya et al., 2025). Chiang et al. (2024b) perform topic-modeling of the prompts submitted to Chatbot Arena. Their top-16 topics include “Poetry Writing Prompts” and “Movie Recommendations and Ratings.” The subjective nature of such topics may make differentiation between top models less meaningful. (3) Ensure higher-quality preference annotations. Arenas could use mediators to perform fine-grained assessments of crowdsourced responses (Don-Yehiya et al., 2025), and categorizing prompts by instruction type (e.g., factual recall, creative generation) to promote more fine-grained model comparisons within categories (Chia et al., 2024).
A complementary line of work on creating high-quality synthetic benchmarks argues that separability—requiring performance gaps between models to be wide enough for leaderboard trends to remain stable under subsampling—should be a main design criterion (Li et al., 2024). At the same time, our findings may suggest that apparent leaderboard differences may be artifacts of noise in the evaluation process rather than genuine performance gaps, which cautions against treating AI leaderboard rankings as definitive indicators of differences in model performance.
@@ A.1 Sensitivity of LLM Arena Rankings Based on Bootstrap Confidence Intervals
Figure 3: Bootstrap-confidence-interval-based rankings on Chatbot Arena (Human Judge).
@@
Figure 4: Bootstrap-confidence-interval-based rankings on Chatbot Arena (LLM Judge).
@@
Figure 5: Bootstrap-confidence-interval-based rankings on Vision Arena.
@@
Figure 6: Bootstrap-confidence-interval-based rankings on Search Arena.
@@
Figure 7: Bootstrap-confidence-interval-based rankings on Webdev Arena.
@@
Figure 8: Bootstrap-confidence-interval-based rankings on MTBench (Human Judge).
@@
Figure 9: Bootstrap-confidence-interval-based rankings on MTBench (LLM Judge).
@@
To construct Table 2, we first recompute the bootstrap-based rankings on the full dataset, apply our method to identify influential preferences, remove those preferences, and then recompute the bootstrap-based rankings. Along with Table 2, we display the plots of the bootstrap-based rankings for the full data and the rankings post-data-dropping in Figures 3, 4, 5, 6, 7, 8 and 9.
@@ A.2 Distinction Between Worst-case Data-dropping Sensitivity and Confidence Intervals
Confidence intervals, such as the bootstrap intervals reported on LMArena (LMArena, 2025), do quantify a form of sensitivity of BT-estimated rankings to variability across samples. However, the sampling-based sensitivity that bootstrap confidence intervals capture is conceptually different from that captured by AMIP. Bootstrap intervals characterize how much an estimate (e.g., the BT score) varies when data are resampled uniformly at random. In contrast, AMIP measures the maximum change in a BT-score difference that can be induced by removing a worst-case small fraction of the data. While frequentist (Gao et al., 2023; Hunter, 2004) confidence intervals methods are meant to capture randomness in the data-generating process, the AMIP targets sensitivity on a single, fixed dataset. This focus on a single sample differs in spirit from the variability across “counterfactual worlds” in that the uncertainty quantification methods are meant to measure. In this sense, the two approaches answer complementary questions about the stability of a sample-based conclusion: the confidence intervals measure sampling uncertainty, while worst-case data-dropping robustness examines whether the ranking is driven by a very small fraction of the observations in the sample. Although Bayesian credible intervals (Leonard, 1977) also operate under the case of a single, fixed dataset, past work has demonstrated that data analyses can be both statistically significant in the Bayesian sense (credible interval does not include zero) and still sensitive to worst-case data dropping (see Bayesian hierarchical model case study in Section 4.4 of (Broderick et al., 2020)). So, analogous to the frequentist case, the AMIP again represents a different and complementary check.
@@ A.3 Uniform Data-Dropping Experiment
To examine the contrast between worst-case data-dropping and dropping random pairs of preferences, we conduct a uniform subsampling experiment. For each arena, we drop $1\%$ of the evaluations uniformly at random, repeat the experiment 100 times, and record the fraction of runs in which the top-ranked model remains unchanged relative to the full arena. For Chatbot Arena (human-judge), we additionally report robustness at a finer scale of $\alpha=0.1\%$.
@@ C.3 Applying AMIP to BT models in logistic form
    Identify the $\left\lfloor\alpha N\right\rfloor$ matchups for which the difference in influence scores is the largest in the negative direction (assuming that player $i$ has a higher estimated BT score than player $j$ on the full data).
@@
    Approximate impact of dropping these $\left\lfloor\alpha N\right\rfloor$ matchups by the sum of the influence score approximations.
@@ Appendix G Verification of AMIP-identified Subsets
In Table 4, we find that all identified $w$-vectors lead to a true change in ranking. And we find that this result holds even when the dropped subset is greater than $\left\lfloor\alpha N\right\rfloor$ of the data (even though the original AMIP makes no claims to an accurate identification of a decision-changing $w$ in this case). However, this does not mean that there are no cases where AMIP fails to surface a vector $w$ that leads to a change in ranking (i.e., false negatives are possible).

### v3

@@ 5.2 Data-dropping Robustness
A growing body of works in statistics and theoretical computer science develops algorithms for assessing whether data analyses are robust to dropping a small, worst-case fraction of the data (Broderick et al., 2020; Kuschnig et al., 2021; Moitra and Rohatgi, 2023; Freund and Hopkins, 2023; Nguyen et al., 2024; Huang et al., 2025a; Rubinstein and Hopkins, 2025). To our knowledge, only one prior work has investigated this question in the context of ranking systems: Shiffman et al. (2023) study the robustness of rankings in gene set enrichment analysis, showing that dropping just a few cells can alter the ranking of p-values derived from the hypergeometric test. In contrast, our work examines ranking robustness in a BT-based ranking system. While Shiffman et al. (2023) analyze p-value rankings, we analyze preference-based rankings of LLMs, extending AMIP (Broderick et al., 2020) to study the robustness of BT-based ranking systems.
@@ 6 Discussion
Sensitivity to worst-case data-dropping is often indicative of low signal-to-noise in the underlying data (Broderick et al., 2020); to help increase signal-to-noise, we recommend three different design-related improvements that AI arenas could make. (1) Collect richer forms of feedback beyond binary preferences (e.g., asking for evaluators’ confidence levels (Méndez et al., 2022). [^6] The weighted logistic regression model used by Chatbot Arena can easily be extended to take in confidence ratings on top of binary preferences. One could imagine implementing this through encoding the confidence rating as a weight in the Win-Counts matrix described in the “Chatbot Arena Leaderboard Calculation (Bradley–Terry model)” Colab notebook. (2) Design more discriminative prompts. Arenas could incorporate a prompt-filtering system to identify and remove uninformative prompts, or create tools to identify prompts requiring specialized knowledge in order to route them to appropriate evaluators (Don-Yehiya et al., 2025). Chiang et al. (2024b) perform topic-modeling of the prompts submitted to Chatbot Arena. Their top-16 topics include “Poetry Writing Prompts” and “Movie Recommendations and Ratings.” The subjective nature of such topics may make differentiation between top models less meaningful. (3) Ensure higher-quality preference annotations. Arenas could use mediators to perform fine-grained assessments of crowdsourced responses (Don-Yehiya et al., 2025), and categorize prompts by instruction type (e.g., factual recall, creative generation) to promote more fine-grained model comparisons within categories (Chia et al., 2024).
A complementary line of work on creating high-quality synthetic benchmarks argues that separability—requiring performance gaps between models to be wide enough for leaderboard trends to remain stable under subsampling—should be a main design criterion (Li et al., 2025). At the same time, our findings may suggest that apparent leaderboard differences may be artifacts of noise in the evaluation process rather than genuine performance gaps, which cautions against treating AI leaderboard rankings as definitive indicators of differences in model performance.
@@ A.1 Sensitivity of LLM Arena Rankings Based on Bootstrap Confidence Intervals
Figure 3: 95% Bootstrap-confidence-interval-based rankings on Chatbot Arena (Human Judge).
@@
Figure 4: 95% Bootstrap-confidence-interval-based rankings on Chatbot Arena (LLM Judge).
@@
Figure 5: 95% Bootstrap-confidence-interval-based rankings on Vision Arena.
@@
Figure 6: 95% Bootstrap-confidence-interval-based rankings on Search Arena.
@@
Figure 7: 95% Bootstrap-confidence-interval-based rankings on Webdev Arena.
@@
Figure 8: 95% Bootstrap-confidence-interval-based rankings on MTBench (Human Judge).
@@
Figure 9: 95% Bootstrap-confidence-interval-based rankings on MTBench (LLM Judge).
@@
To construct Table 2, we first compute the bootstrap-based rankings on the full dataset, apply our method to identify influential preferences, remove those preferences, and then recompute the bootstrap-based rankings. Along with Table 2, we display the plots of the bootstrap-based rankings for the full data and the rankings post-data-dropping in Figures 3, 4, 5, 6, 7, 8 and 9.
@@ A.2 Distinction Between Worst-case Data-dropping Sensitivity and Confidence Intervals
Confidence intervals, such as the bootstrap intervals reported on LMArena (LMArena, 2025), do quantify a form of sensitivity of BT-estimated rankings to variability across samples. However, the sampling-based sensitivity that bootstrap confidence intervals capture is conceptually different from that captured by AMIP. Bootstrap intervals characterize how much an estimate (e.g., the BT score) varies when data are resampled uniformly at random. In contrast, AMIP measures the maximum change in a BT-score difference that can be induced by removing a worst-case small fraction of the data. While frequentist (Gao et al., 2023; Hunter, 2004) confidence interval methods are meant to capture randomness in the data-generating process, the AMIP targets sensitivity of a single, fixed dataset. This focus on a single sample differs in spirit from the variability across “counterfactual worlds” that uncertainty quantification methods are meant to measure. In this sense, the two approaches answer complementary questions about the stability of a sample-based conclusion: the confidence intervals measure sampling uncertainty, while worst-case data-dropping robustness examines whether the conclusion is driven by a very small fraction of the observations in the sample. Although Bayesian credible intervals (Leonard, 1977) also operate under the case of a single, fixed dataset, past work has demonstrated that data analyses can be both statistically significant in the Bayesian sense (credible interval does not include zero) and still sensitive to worst-case data dropping (see Bayesian hierarchical model case study in Section 4.4 of (Broderick et al., 2020)). So, analogous to the frequentist case, the AMIP again represents a different and complementary check.
@@ A.3 Uniform Data-Dropping Experiment
To examine the contrast between worst-case data-dropping and dropping random pairwise preferences, we conduct a uniform subsampling experiment. For each arena, we drop $1\%$ of the evaluations uniformly at random, repeat the experiment 100 times, and record the fraction of runs in which the top-ranked model remains unchanged relative to the full arena. For Chatbot Arena (human-judge), we additionally report robustness at a finer scale of $\alpha=0.1\%$.
@@ C.3 Applying AMIP to BT models in logistic form
    Identify the $\lfloor\alpha N\rfloor$ matchups for which the difference in influence scores is the largest in the negative direction (assuming that player $i$ has a higher estimated BT score than player $j$ on the full data).
@@
    Approximate impact of dropping these $\lfloor\alpha N\rfloor$ matchups by the sum of the influence score approximations.
@@ Appendix G Verification of AMIP-identified Subsets
In Table 4, we find that all identified $w$-vectors lead to a true change in ranking. And we find that this result holds even when the dropped subset is greater than $\lfloor\alpha N\rfloor$ of the data (even though the original AMIP makes no claims to an accurate identification of a decision-changing $w$ in this case). However, this does not mean that there are no cases where AMIP fails to surface a vector $w$ that leads to a change in ranking (i.e., false negatives are possible).